On a particular test, IQ scores for adults follow a bell-shaped distribution with mean= 100 and a variance of 4. In a random sample of 300 adults, approximately how many would have IQs greater than 104?
Answer by stanbon(75887)

On a particular test, IQ scores for adults follow a bell-shaped distribution with mean= 100 and a variance of 4. In a random sample of 300 adults, approximately how many would have IQs greater than 104?
----------
z(104)= (104-100)/4 = 1
---
P(x > 104) = P(z > 1) = normalcdf(1,100) = 0.1587
==================
Cheers,
Stan H.
==================